OpenStudy (cutiecomittee123):

how do i solve this system of equations (x+1)^2/25 + (y+2)^2/4 =1 y=(x+1)^2+1 I know that the first one is an ellipse and the second one is a parabola therefore there can be up to 3 intersection points.

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

i guess you can replace \(y\) in the first equation by \((x+1)^2+1\) but maybe there is an easier way

OpenStudy (anonymous):

want to try that? (btw they do not intersect)

OpenStudy (cutiecomittee123):

its a hard one

OpenStudy (cutiecomittee123):

the two graphs dont intersect?

OpenStudy (anonymous):

no they do not

OpenStudy (anonymous):

damn that is wrong

OpenStudy (anonymous):

thinking...

OpenStudy (anonymous):

ok how about this :\[y=(x+1)^2+1\] so \[(x+1)^2=y-1\]

OpenStudy (anonymous):

replace \((x+1)^2\) in the ellipse with \(y-1\) you will get a quadratic equation in \(y\)

OpenStudy (anonymous):

@iambatman how does that seem to you?

OpenStudy (anonymous):

\[\frac{y-1}{25}+\frac{(y+2)^2}{4}=1\]lets try that

OpenStudy (anonymous):

\[4(y-1)+25(y+2)^2=100\] thats better

OpenStudy (anonymous):

i must be screwing up somehow

OpenStudy (cutiecomittee123):

omg i got distracted

OpenStudy (anonymous):

i am still messing up i have to figure this out

OpenStudy (cutiecomittee123):

im really confused with it

OpenStudy (anonymous):

me too but i will figure it out somehow

OpenStudy (cutiecomittee123):

I am just going to graph them both and see what happens

OpenStudy (anonymous):

i did that first

OpenStudy (cutiecomittee123):

https://www.desmos.com/calculator

OpenStudy (cutiecomittee123):

Yeah they don't intersect what so ever. Well anyways that is fine. the question just says to graph the system of equations and that's exactly what we both did, thanks:) lol sorry for the confusion.

OpenStudy (anonymous):

oooh! it said graph...